I have two questions that I need help on.
The first one is: The length of a rectangle is twice the width. the area is 50 yd^2. fin the length and the width.
The second one is: A student opens a mathematics book to two facing pages. the product of the page numbers is 2162. Find the page numbers.
Can someone please help me with these two? I would really appreciate it, and I thank you in advance for your help.

Answer by user_dude2008(1862)

"The length of a rectangle is twice the width. the area is 50 yd^2" -----> LW=50
and L=2W

--->
(2W)W=50

2W^2=50
W^2=25

W=5...

L=2*5=10

Ans: width: 5, length: 10


.................................


" A student opens a mathematics book to two facing pages" ----> x(x+1)=2162

x^2+x=2162

x^2+x-2162=0

(x-46)(x+47)=0

x=46, x=-47

Ans: pages are 46 and 47